(x+6)(x-19)(x+1)=0 what are the x's???
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

(x+6)(x-19)(x+1)=0 what are the x's???
====
Since the product is zero, one of those factors must be zero.
----
x+6 = 0 or x-19 = 0 or x+1 = 0
x = -6 or x = 19 or x = -1
===============================
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi
The factor theorem states that a polynomial f(x) has a factor (x − k) if and only if f(k) = 0
f(x) = (x+6)(x-19)(x+1)=0 |As Stated in the Question, therefore:
(x+6)= 0
x = -6
(x-19)= 0
x = 19
(x+1)=0
x = -1
Solution Set is: -6,-1,19